7.

Annett, starting from a point A, walks 8km on bearing of 2650


to a point B. From B she walks on


bearing of 1350


for 9km to a point C. Find the bearing of A from C.


1
Expert's answer
2021-09-03T15:14:39-0400

The bearing of a point is the number of degrees in the angle measured in a clockwise direction from the north line to the line joining the centre of the compass with the point. 


"\overrightarrow{AB}=(8\sin(265\degree), 8\cos(265\degree))"

"\overrightarrow{BC}=(9\sin(135\degree), 9\cos(135\degree))"

"\overrightarrow{AC}=\overrightarrow{AB}+\overrightarrow{BC}"

"=(8\sin(265\degree)+9\sin(135\degree), 8\cos(265\degree)+9\cos(135\degree))"

"\tan \theta=\dfrac{8\cos(265\degree)+9\cos(135\degree)}{8\sin(265\degree)+9\sin(135\degree)}"

The bearing of A from C


"90\degree-\theta=90\degree-\tan^{-1}(\dfrac{8\cos(265\degree)+9\cos(135\degree)}{8\sin(265\degree)+9\sin(135\degree)})"

"\approx12.81\degree"